2.7a PMEL-EDD Accomplishments on RHB Vents Leg II - Chris Meinig

(These are in no particular order)

1. Extensive work was done in port to remove a Trackpoint II transducer and install and test an Edgetech PS-8000 transducer on the extensible mast of the RHB. The PS-8000 rangemeter electronics was installed in the CME office and cables were run to the ROV control room.

2. The NemoNet 2000 camera was tested, deployed and positioned with ROPOS near Marker 33. A temperature tripod with two RTD was place in the clump of tubeworms. Nominal temperature in the worms was 11 deg C. Several tiles and temperatures were taken to make sure camera was operating properly.

3. The NemoNet surface buoy was deployed about 300m due west of the camera. Buoy electronics was monitored while the camera was interrogated from the RHB. All tests checked out ok. Anchor position was surveyed from the surface.

4. The NemoNet 1999 camera was recovered using a dropped mooring with a line canister. The 1999 camera was moved near the mooring and made fast by ROPOS with a snap hook. The acoustic release was tripped from the surface and the entire mooring was successfully recovered.

5. Eleven Benchmark Acoustic Extensometers were deployed to the seafloor in two elevator drops. Each elevator drop had to be surveyed in from the surface to provide a dive target. The extensometers were then place in steel benchmarks already on the seafloor. Instruments#2-9 were interrogated via infrared link and the data was downloaded to a computer.

6. 4 "classic" extensometers were deployed via the elevator and placed by ROPOS at predetermined positions. Elevator was successfully recovered.

7. The hot fluid sampler was installed on three ROPOS dives. Several minor problems with telemetry and power were troubleshot and solved.

8. 5 mini-benchmarks were deployed on two separate ROV dives. The cage was rigged to drop the mini-benchmark from an actuated cylinder.

9. An MTR mooring was surveyed from the surface using Workboat and found to have the wrong acoustic release. #49 & #50 had been switched.

10. Two flooded Hobo temperature probes were diagnosed for data recovery. No data was recovered. Instruments should be sent to the manufacturer.

11. Three TR-6000 acoustic beacons were deployed and recovered.

Please give special thanks to CME Mike Gowan and the ship's engineers for helping to repair several instrument brackets on short notice and giving us access to their shop and tools.